# Systematic study of confinement induced effects on atomic electronic structure

### v1

@@ 5 Summary and conclusion
To the best of our knowledge, a robust and systematic study of atoms in confinement has not been hitherto published in the literature. Such a study both needs to use a suitable numerical method and to consider the changing nature of the ground state as a function of the confinement, and we did not find a single publication that satisfies both these criteria, while considering an appreciable portion of the periodic table instead of just some specific subrows or groups.
We therefore took a fully numerical approach based on the finite element method (FEM) to study atoms in hard-wall confinement. We chose to represent the atoms with spherically averaged, spin-restricted or spin-polarized densities, and carried out calculations with the Perdew–Burke–Ernzerhof (PBE) exchange-correlation functional.22, 23 We compared the ionization energies of unconfined atoms predicted by spin-restricted or spin-polarized calculations to experiment, and observed the latter to be unsurprisingly more accurate, while even the spin-restricted calculations succeed in predicting periodic trends.
We then carried out a systematic study of the total energies of the ground and low lying excited states of the H–Xe atoms as well as their monocations as a function of the confinement radius $r_{c}$, paying special focus on how the ground state evolves as a function of confinement. We demonstrated the significance of considering confinement induced changes in the electronic structure by comparing ionization radii computed for fixed electronic configurations to those obtained in calculations, where the electronic configuration can relax as a function of the confinement radius. Depending on the atom, the relaxation can result in either an increase or a decrease of the ionization radius, and the largest differences are seen for the pre-$d$ and early $d$ elements.
We summarize our analysis of the calculations as follows.
@@
  • Valence $s$ electrons are highly unfavoured under strong confinement: already the Li and Be atoms show an $ns\to np$ electron transfer.
@@
Our systematic examination of the low lying configurations of the elements H–Xe highlights drastic changes in the electronic structure of confined atoms and provides a base for future studies with higher-accuracy wave function methods.
Like most studies on confined atoms, the present study was limited to the use of integer occupation numbers for the $s$, $p$, $d$, and $f$ shells to maximize physical interpretability. However, as we discussed in section 4.2.2, the use of non-integral occupation numbers could be useful for transition metals, for instance,46, 47, 28 and we hope to revisit such calculations with state-of-the-art optimization algorithms such as optimal damping57 in future work.
Our study was limited to the H–Xe atoms, because relativistic effects are well known to be essential for heavy elements.58, 59 Relativistic calculations of atoms in confinement should be feasible using the same approach, and such calculations might be carried out with the recent implementation of Čertík et al.60, for example.
@@
H.Å. thanks the Finnish Society for Sciences and Letters for funding. S.L. thanks the Academy of Finland for financial support under project numbers 350282 and 353749.
@@
The SI PDF file contains the following data. Plots of the total energy of the low lying states for the H–Xe atoms and their monocations as a function of the confinement radius. Detailed analysis of the evolution of the ground state, and state crossings with the ground state configuration of the unconfined atom. Estimated atomic sizes and excitation energies relative to the ground state for the studied configurations in the unconfined atom. Values for ionization radii via Janak’s theorem for both the fixed and relaxed configuration approaches. All of the above sets of results are given both for the spin-restricted and spin-polarized formalisms.
The raw energies for all calculations are included in the comma separated values (CSV) format in a compressed archive.

### v2

@@ 5 Summary and conclusion
To the best of our knowledge, a robust and systematic study of atoms in confinement has not been hitherto published in the literature. Such a study both needs to use a suitable numerical method and to consider the changing nature of the ground state as a function of the confinement. We did not find a single publication that satisfies both these criteria while also considering an appreciable portion of the periodic table instead of just some specific subrows or groups.
We therefore took a fully numerical approach based on the finite element method (FEM) to study atoms in hard-wall confinement. We chose to represent the atoms with spherically averaged, spin-restricted or spin-polarized densities, and carried out calculations at three levels of density functional approximations: the local density approximation employing the Perdew–Wang (PW92) correlation functional, the generalized-gradient approximation (GGA) employing the Perdew–Burke–Ernzerhof (PBE) exchange-correlation functional, and the meta-GGA approximation employing the r2SCAN exchange-correlation functional. We compared the ionization energies of unconfined atoms predicted by spin-restricted or spin-polarized calculations to experiment, and unsurprisingly observed the latter to be more accurate, while even the spin-restricted calculations succeed in predicting periodic trends.
We demonstrated the significance of considering confinement induced changes in the electronic structure by comparing ionization radii computed for fixed electron configurations to those obtained in calculations in which the electron configuration can relax as a function of the confinement radius. Depending on the atom, this relaxation can result in either an increase or a decrease of the ionization radius due to competing effects in the neutral atom and its cation, and the largest differences are seen for the pre-$d$ and early $d$ elements.
As a side result of our study, we also examined various atomic size estimates. We found severe issues in the van der Waals (vdW) radius estimate proposed by Bader et al.51: it is not always defined, it has severe issues with numerical stability even with huge numerical basis sets, and its predictions do not agree with basic physical insights for the $1s^{2}$ ground and $1s^{1}2p^{1}$ and $1s^{1}3d^{1}$ excited configurations of the He atom. We proposed a related estimate which is always defined, does not suffer from severe numerical problems like the aforementioned estimate, and also predicts the aforementioned states of the He atom to be increasingly diffuse in character, in agreement with the covalent radius estimate of Slater50. We expect our proposed vdW radius estimate to be useful in future work.
We then carried out a systematic study of the total energies of the ground and low lying excited states of the H–Xe atoms as well as their monocations as a function of the confinement radius $r_{c}$, paying special focus on how the ground state evolves as a function of confinement. We have confirmed electron shifts made in previous studies on confined atoms, but also made novel observations. We summarize our analysis of the calculations as follows.
@@
  • Valence $s$ electrons are highly unfavoured under strong confinement: already the Li and Be atoms show a $2s\to 2p$ electron shift.
@@
Importantly, we saw that the dissimilar density functional approximations were in excellent agreement in calculations of ionization radii, and that the qualitative behavior of the confined atoms is similar for dissimilar density functionals. We are therefore confident that our results are not artefacts of the employed level of theory, and that similar results would be obtained even with high-level ab initio calculations.
Like most studies on confined atoms, the present study was limited to the use of integer occupation numbers for the $s$, $p$, $d$, and $f$ shells to maximize physical interpretability. However, as we discussed in section 4.2.2, the use of non-integral occupation numbers could be useful for transition metals, for instance,47, 48, 37 and we hope to revisit such calculations with state-of-the-art optimization algorithms such as optimal damping83 in future work.
Our study was limited to the H–Xe atoms, because relativistic effects are well known to be essential for heavy elements.84, 85 Relativistic calculations of atoms in confinement should be feasible using the same approach, and such calculations might be carried out with the recent implementation of Čertík et al.86, for example.
5.1 Connections to basis set design
It is interesting to draw a parallel here to basis set design, since the application of confinement potentials in the construction of numerical atomic orbital (NAO) basis sets as originally proposed by by Averill and Ellis87 was the original motivation of our work. While Averill and Ellis87 employed a finite potential barrier, Sankey and Niklewski88 proposed building NAO basis sets using a hard-wall potential to enforce the locality of the resulting basis set. In later works, Sankey and coworkers refer to the arising orbitals as “fireballs” due to their excited nature relative to the unconfined atom.89, 90 The technique of Sankey and Niklewski88 has been later used by many other authors as well, such as Sánchez-Portal et al.91, Basanta et al.92, Kenny and Horsfield93, and Nakata et al.94.
The demands for various chemical elements’ atomic orbital basis sets for electronic structure calculations95, 96, 97 are inherently tied to the elements’ chemistry. Importantly, the relevance of distinct angular momentum $l$ orbitals for atoms is expected to be a continuous function: as we move to heavier elements, we first see an increasing importance of $p$ functions, followed by increasing importance of $d$ functions, $f$ functions and so on. Unlike the occupied orbitals in an SCF calculation on an atom, the importance of polarization functions in a basis set is not a step function for which we can certainly say whether a specific element requires $s$, $p$, $d$, $f$, etc. functions, or not.
This continuity is a challenge when predicting the relevance of distinct orbitals to chemical bonding. The advantage of our arguably simple model is that we are able to highlight these chemical characteristics of the studied atoms, that is, the relative importance of the electron configurations as a function of confinement. The physicality of the model is therefore obvious. For example, we saw in this work that the simple hard-wall confinement model predicts the importance of the $3d$ orbital in K and Ca, which is obvious from the large transition radius $r_{c}\approx 4.5a_{0}$.
[…]
Martin104 found a physical interpretation for this effect: the $3d$ Rydberg orbital on the chlorine atom(s) can accept back-bonding contributions from the oxygen atoms. Mehta and Martin105 also point out that the $3d$ orbital comes down in energy when the atom becomes more positively charged. Indeed, atomic orbitals determined for cations are commonly used to introduce additional flexibility in NAO basis sets.106 Mehta and Martin105 also made the discovery that the fourth-row $5p$ block analogously exhibits a heightened importance of $4f$ Rydberg orbitals.
Based on the above findings, we expect that heavier elements will exhibit electron shifts to the $4f$ orbital in weaker and weaker confinement, since the $4f$ orbital will come down in energy as the nuclear charge is increased.
@@ 5.1 Connections to basis set design
We thank professor Pekka Pyykkö for comments on the manuscript, and Monica Lindholm for designing the cover graphics for the manuscript. H.Å. thanks the Finnish Society for Sciences and Letters for funding. S.L. thanks the Academy of Finland for financial support under project numbers 350282 and 353749.
@@
The SI PDF file contains the following data. Comparison of the ionization energies with the PW92 and r2SCAN functional. Values of the ionization radii via Janak’s theorem for both the fixed and relaxed configuration approaches. Detailed analysis of the evolution of the ground state of the H–Xe atoms, and state crossings with the ground state configuration of the unconfined atom. Estimated atomic sizes and excitation energies relative to the ground state for the studied configurations in the unconfined atom. Plots of the total energy of the low lying states of the H–Xe atoms and their monocations as a function of the confinement radius. All of the above sets of results are given both for the spin-restricted and spin-polarized formalisms for all the functionals considered in this work: PW92, PBE, and r2SCAN. The raw energies for all calculations are included in comma separated values (CSV) format in a compressed archive.